What is the forecast for July using a four-month moving average? DEC 96 APR 46 JAN NIOOW 39 40 FEB 42 MAR 48
Answer #1

Forecast for month n = (An-1 + An-2 + An-3 + An-4)/4

FMay = (46+48+42+40)/4 = 44

FJune = (44+46+48+42)/4 = 45

FJuly = (45+44+46+48)/4 = 45.75